OpenStudy (marigirl):

check my answer please: When three fair coins are tossed, calculate the conditional probability that two heads and a tail turn up, given that first coin came up heads. P(A) is that 2 heads and 1 tail P(B) is 1 one is heads. since we are considering that first one is going to be heads, then P(a) is 1 head and 1 tail? so (1/4) / (1/2)

jimthompson5910 (jim_thompson5910):

` given that first coin came up heads.` so we don't have to factor this in. We just need to consider the other two coins

jimthompson5910 (jim_thompson5910):

P(second one is heads) = 1/2 P(third one is tails) = 1/2 P(second one is heads AND third one is tails) = P(second one is heads)*P(third one is tails) P(second one is heads AND third one is tails) = (1/2)*(1/2) P(second one is heads AND third one is tails) = 1/4

OpenStudy (marigirl):

okay.. and if we do that means we are calculating the event happening twice which is not what we want .. yea?

OpenStudy (marigirl):

thanks!